# LEGO-Puzzles: How Good Are MLLMs at Multi-Step Spatial Reasoning?

### v1

@@ Abstract
Multi-step spatial reasoning entails understanding and reasoning about spatial relationships across multiple sequential steps, which is crucial for tackling complex real-world applications, such as robotic manipulation, autonomous navigation, and automated assembly. To assess how well current Multimodal Large Language Models (MLLMs) have acquired this fundamental capability, we introduce LEGO-Puzzles, a scalable benchmark designed to evaluate both spatial understanding and sequential reasoning in MLLMs through LEGO-based tasks. LEGO-Puzzles consists of 1,100 carefully curated visual question-answering (VQA) samples spanning 11 distinct tasks, ranging from basic spatial understanding to complex multi-step reasoning. Based on LEGO-Puzzles, we conduct a comprehensive evaluation of state-of-the-art MLLMs and uncover significant limitations in their spatial reasoning capabilities: even the most powerful MLLMs can answer only about half of the test cases, whereas human participants achieve over 90% accuracy. In addition to VQA tasks, we evaluate MLLMs’ abilities to generate LEGO images following assembly illustrations. Our experiments show that only Gemini-2.0-Flash and GPT-4o exhibit a limited ability to follow these instructions, while other MLLMs either replicate the input image or generate completely irrelevant outputs. Overall, LEGO-Puzzles exposes critical deficiencies in existing MLLMs’ spatial understanding and sequential reasoning capabilities, and underscores the need for further advancements in multimodal spatial reasoning.
@@ 1 Introduction
Research on evaluating MLLMs’ multi-step spatial reasoning capabilities remains largely unexplored. Existing studies primarily focus on assessing the spatial understanding capability, which pertains to the comprehension of a static scene. Some works [21, 30, 53] employ synthetic environments to render multiple simple 3D objects and then query the spatial relationships between them. However, such question-answering (QA) tasks tend to be overly simplistic for MLLMs to handle, lacking the diversity and complexity of real-world scenarios. Other studies [38, 31] construct spatial understanding tasks based on natural images, but this approach often involves manual annotations, which may limit scalability. Moreover, most existing evaluations rarely evaluate reasoning over sequences of spatial transformations or actions, leaving the multi-step aspect of spatial reasoning largely unaddressed.
@@
In this work, we take inspiration from a common recreational activity, LEGO construction, to design a comprehensive evaluation framework for assessing the multi-step spatial reasoning capabilities of MLLMs. The assembly process of a complete LEGO model typically encompasses dozens or even hundreds of discrete construction steps, providing an ideal foundation for testing sequential reasoning abilities. Each step requires accurate comprehension of geometry, orientation, and connection mechanisms of LEGO pieces to successfully follow the provided illustrations. Based on publicly available LEGO projects with detailed step-by-step assembly instructions, we introduce LEGO-Puzzles, a novel benchmark specifically engineered to evaluate MLLMs’ multi-step spatial reasoning capabilities. In total, LEGO-Puzzles encompasses a diverse collection of over 1,100 carefully curated visual question-answering (VQA) pairs spanning 11 distinct tasks, which fall into three major categories. First, we develop a set of fundamental tests to assess MLLMs’ basic spatial understanding capabilities, including recognition of height relationships, rotational transformations, adjacency patterns, and viewpoints within 3D space. Building upon this foundation, we construct both single-step and multi-step sequential reasoning evaluations based on LEGO assembly sequences to examine models’ sequential reasoning ability. These advanced tests include identifying the configuration of intermediate assembly states (single-step) or determining the correct order of multiple intermediate LEGO states (multi-step).
LEGO-Puzzles offers several distinctive advantages compared to existing spatial understanding benchmarks: 1) Enhanced visual richness. Unlike synthetic datasets such as CLEVR [21, 30], which utilize rendered primitive shapes, LEGO-based questions present significantly greater visual complexity and diversity. 2) Superior scalability. A single LEGO assembly instruction manual can generate hundreds of unique evaluation questions, which enables efficient benchmark expansion with minimal additional resource investment.
@@
Leveraging LEGO-Puzzles, we conduct comprehensive evaluations of 20 state-of-the-art MLLMs, including proprietary models such as GPT-4o and Gemini-2.0-Flash, as well as leading open-source alternatives [8, 50, 56, 43]. Our experimental results reveal a substantial gap between current MLLMs and human-level proficiency. Even the strongest models struggle with basic spatial understanding tasks, such as accurately identifying the height of LEGO pieces and determining adjacency relationships in 3D space. Among open-source models, only a few achieve performance notably above random guessing across different tasks.
@@
Beyond VQA tasks, LEGO-Puzzles also enables the assessment of spatially grounded image generation. For instance, given an assembly illustration, an MLLM is tasked with generating an image of the intermediate state following the specified assembly operation. In these generation tests, most of the evaluated models fail completely, either disregarding the provided instructions or generating images that are entirely irrelevant to the intended LEGO configuration.
@@
In summary, our novel benchmark LEGO-Puzzles provides a comprehensive evaluation of the spatial understanding and sequential reasoning capabilities of MLLMs. Our main contributions are as follows:
  • A novel benchmark for spatial understanding. Based on LEGO constructions, our benchmark LEGO-Puzzles offers natural and diverse test cases for evaluating the spatial understanding capabilities of MLLMs, with improved visual richness and scalability over existing datasets.
@@
  • Evaluation of multi-step spatial reasoning. Built upon LEGO’s step-by-step building process, LEGO-Puzzles is the first benchmark explicitly designed to assess multi-step spatial reasoning, where each task requires reasoning over up to $7$ LEGO construction steps.
  • Comprehensive assessment on visual question answering and image generation. LEGO-Puzzles assesses the spatial reasoning capability of LLMs across both VQA and image generation tasks, providing a comprehensive assessment of their ability to comprehend and process spatial information in a human-like manner.
@@ 3 LEGO-Puzzles
In this section, we introduce LEGO-Puzzles, a diverse and comprehensive benchmark designed to evaluate the multi-step spatial reasoning capability of MLLMs in detail. Specifically, we first introduce the motivation and definition of each task in Section 3.1. Then, we introduce our dataset curation process, including data collection, question-answer generation, and quality control, in Section 3.2.
@@ 3.1 Task Definition
To enhance the evaluation of multi-step spatial reasoning for Multimodal Large Language Models (MLLMs), we define three primary categories of tasks based on insights from cognitive psychology and human experiences in developing relevant skills [40, 5, 55]. Using LEGO building as a concrete example of how humans develop spatial intelligence, we find that individuals typically engage in the following processes: First, they must understand the spatial relationships between each LEGO piece and how these pieces relate from different perspectives in 3D space. Next, they need to reason through the dependencies and assembly logic of each block at every step of the building process. Finally, they extend their reasoning to multi-step reasoning across the entire assembly sequence. To achieve this, our tasks range from fundamental spatial understanding (36.4%) to single-step sequential reasoning (36.4%) and, ultimately, to multi-step sequential reasoning (27.3%), as illustrated in Figure 1. Below, we provide further details on each task.
Task 1: Spatial Understanding. (1) Height: Distinguish the relative heights of LEGO objects. (2) Adjacency: Determine whether LEGO objects are adjacent or separated. (3) Rotation: Calculate the angle of rotation between a LEGO object and its corresponding rotated version. (4) Multiview: Predict the current LEGO status from different viewpoints. Task 2: Single-step Sequential Reasoning. (5) Rotation Status: Assess the rotation status of the next LEGO pieces during assembly. (6) Position: Identify the correct assembly position for the next LEGO pieces. (7) Next-Step: Determine the next LEGO status based on the current status and the upcoming pieces. (8) Dependency: Identify the necessary LEGO pieces required to transition from the current to the next status. Task 3: Multi-Step Sequential Reasoning. (9) Backwards: Identify the correct LEGO status in the assembly pipeline of the LEGO object. (10) Ordering: Predict the correct assembly order of the provided final LEGO images. (11) Outlier: Detect the LEGO status that does not belong to the provided assembly sequence.
Figure 1: Problem Statistics in LEGO-Puzzles.
In conclusion, LEGO-Puzzles consists of over 1,100 visual question-answering (VQA) pairs derived from 407 LEGO building instructions, encompassing 11 tasks across spatial understanding, single-step and multi-step sequential reasoning. In addition to VQA tasks, We further extend several sub-tasks of spatial understanding (Rotation* and Multiview*) and single-step sequential reasoning (Position*, Dependency* and Next-Step*) to include image generation following [57], as part of the visual Generation evaluation of MLLMs.
@@ 3.2 Dataset Curation
Figure 3: Data curation pipeline. Our pipeline first collects a diverse set of LEGO building instructions to render and extract LEGO images in a unified format. Next, we generate question-answer pairs by using a combination of human annotation and predefined question templates. Finally, we implement three quality control strategies to ensure the accuracy, consistency, and reliability of the data.
@@
As illustrated in Figure 3, our pipeline consists of three key steps: data collection, question-answer generation, and quality control. This design ensures the scalability, accuracy, and reliability of our data.
Data Collection. Data collection consists of three stages. First, we collect a diverse set of open-source LEGO source files from the Internet, which include comprehensive step-by-step LEGO building instructions, visualizations, and the required LEGO pieces for each step. Notably, the camera perspective remains consistent across all steps within a specific instruction, ensuring temporal, spatial, and logical coherence throughout the building process. Second, we render LEGO building instruction files as PDF files using a publicly available rendering software, Studio [^1] https://www.bricklink.com/v3/studio/download.page. This tool enables us to adjust default rendering settings to construct tasks that evaluate spatial relationships at varying levels of complexity. Specifically, for the Rotation and Multiview tasks, we utilize the Persistence of Vision Raytracer (POV-Ray) style and modify the lighting strength to generate realistic LEGO images from different angles. For the Backward task, we also edit attributes such as color, quantity, and assembly positions of pieces to create erroneous images. Finally, we use PDF-Extract [^2] https://github.com/opendatalab/PDF-Extract-Kit to extract all LEGO pieces and objects of interest from the rendered PDF files. All images are systematically organized according to a unified naming standard and prepared for question-answer generation across different tasks.
Question-Answer Generation. To ensure the scalability of our pipeline, we design several task-specific templates for question-answer generation. See Figure 4 for an example. Each data example includes an instruction, a question, and an answer. To meet the requirements of different tasks, we create LEGO sequences of varying lengths. Note that the image token <image x> serves as a placeholder for the corresponding image input here.
Figure 4: Task-specific template. Our question-answer template includes instructions, questions, and answers. Here, we provide an example from the Position task for reference.
Quality Control. We implement a rigorous human review process to maintain high quality and minimize errors. Specifically, we carefully examine the consistency between LEGO objects in source files and rendered PDF files, conducting checks for duplication, adherence to standards, and formatting in the generated images. Additionally, we apply cross-validation to ensure that each question-answer pair aligns with its task-specific template and that the image annotations are accurate, as verified by multiple annotators. This rigorous process ensures high-quality and reliable evaluation.
4 Evaluation on LEGO-Puzzles
@@ 4.1 Experimental Setting
Benchmark Models. We extensively evaluate 20 models, covering a diverse range of architectures, sizes, and training processes for Spatial Understanding and Sequential Reasoning tasks. For open-source models, we evaluate MiniCPM-V2.6 [59], Qwen2-VL-[7B/72B] [50], Qwen2.5-VL-[7B/72B] [4], InternVL2.5-[8B/78B] [7], VILA1.5-13B [32], Idefics3-8B [24], DeepSeek-VL2-[Tiny/Small] [56], Pixtral-12B [1], LLaVA-OneVision-7B [26], and EMU3 [54]. For proprietary models, we evaluate Claude-3.5-Sonnet [2], Gemini-1.5-Flash, Gemini-1.5-Pro, Gemini-2.0-Flash [48], GPT-4o (20241120), and GPT-4o-mini [42]. For the additional image Generation evaluation, we evaluate the open-source models Emu2 [47], GILL [23], and Anole [9], as well as the proprietary models GPT-4o and Gemini-2.0-Flash, all of which support long-range sequence input and image output. Moreover, all evaluations are conducted in a zero-shot setting for a fair comparison.
@@
Evaluation Metrics. We calculate the accuracy (%) of verbal answers to multiple-choice questions using exact matching as our primary metric, following Duan et al. [11]. When models fail to generate an answer in the required format, we utilize the ChatGPT-0125 [13] method from VLMEvalKit [11] as a fallback option. Additionally, we randomly select 20 questions from each task to create LEGO-Puzzles-Lite, resulting in a total of 220 question-answer pairs. This dataset is designed to investigate the performance gap between human intelligence, as evaluated by additional human annotators, and current models in patial Understanding and Sequential Reasoning. In the Generation evaluation, traditional metrics such as FID [20], CLIPScore [19, 15], and X-IQE [6] are inadequate for assessing interleaved outputs in visual answers. Therefore, we enlist human experts to evaluate performance based on appearance similarity and instruction following, using a scoring scale from 0 to 3. This approach is necessary due to biases present in VLM-based scoring [35].
@@
Table 1: Full Evaluation Results of 18 MLLMs on LEGO-Puzzles. Dark Gray indicates the best performance for each task among all models and Light Gray indicates the best result among open-source model. We also highlight the top three models based on their overall performance, using Dark Green, Medium Green, and Light Green, respectively.
@@
Table 2: Comparing Top-Performing MLLMs with Human Proficiency on LEGO-Puzzles-Lite. The best results are marked in bold. The top three overall performances are highlighted in Dark Green, Medium Green, and Light Green, respectively.

### v2

@@ Abstract
Multi-step spatial reasoning entails understanding and reasoning about spatial relationships across multiple sequential steps, which is crucial for tackling complex real-world applications, such as robotic manipulation, autonomous navigation, and automated assembly. To assess how well current Multimodal Large Language Models (MLLMs) have acquired this fundamental capability, we introduce LEGO-Puzzles, a scalable benchmark designed to evaluate both spatial understanding and sequential reasoning in MLLMs through LEGO-based tasks. LEGO-Puzzles consists of 1,100 carefully curated visual question-answering (VQA) samples spanning 11 distinct tasks, ranging from basic spatial understanding to complex multi-step reasoning. Based on LEGO-Puzzles, we conduct a comprehensive evaluation of 20 state-of-the-art MLLMs and uncover significant limitations in their spatial reasoning capabilities: even the most powerful MLLMs can answer only about half of the test cases, whereas human participants achieve over 90% accuracy. Furthermore, based on LEGO-Puzzles, we design generation tasks to investigate whether MLLMs can transfer their spatial understanding and reasoning abilities to image generation. Our experiments show that only GPT-4o and Gemini-2.0-Flash exhibit a limited ability to follow these instructions, while other MLLMs either replicate the input image or generate completely irrelevant outputs. Overall, LEGO-Puzzles exposes critical deficiencies in existing MLLMs’ spatial understanding and sequential reasoning capabilities, and underscores the need for further advancements in multimodal spatial reasoning.
@@ 1 Introduction
Research on evaluating MLLMs’ multi-step spatial reasoning capabilities remains largely unexplored. Existing studies primarily focus on assessing the spatial understanding capability, which pertains to the comprehension of a static scene. Some works [20, 29, 52] employ synthetic environments to render multiple simple 3D objects and then query the spatial relationships between them. However, such tasks tend to be overly simplistic for MLLMs to handle, lacking the diversity and complexity of real-world scenarios. Other studies [37, 30] construct spatial understanding tasks based on natural images, but this approach often involves manual annotations, which may limit scalability. Moreover, most existing evaluations rarely evaluate reasoning over sequences of spatial transformations or actions, leaving the multi-step aspect of spatial reasoning largely unaddressed.
@@
In this work, we take inspiration from a common recreational activity, LEGO construction, to design a comprehensive evaluation framework for assessing the multi-step spatial reasoning capabilities of MLLMs. The assembly process of a complete LEGO model typically encompasses dozens or even hundreds of discrete construction steps, providing an ideal foundation for testing sequential reasoning abilities. Each step requires accurate comprehension of geometry, orientation, and connection mechanisms of LEGO pieces to successfully follow the provided illustrations. Based on publicly available LEGO projects with detailed step-by-step assembly instructions, we introduce LEGO-Puzzles, a benchmark specifically engineered to evaluate MLLMs’ multi-step spatial reasoning capabilities. In total, LEGO-Puzzles encompasses a diverse collection of over 1,100 carefully curated visual question-answering (VQA) pairs spanning 11 distinct tasks, grouped into three categories (Fig. 1): a set of fundamental tests to assess MLLMs’ basic spatial understanding capabilities, including recognition of height relationships, rotation angles, 3D viewpoints, etc. Building upon this foundation, we construct both single-step and multi-step sequential reasoning evaluations based on LEGO assembly sequences to examine models’ sequential reasoning ability. In addition, a single LEGO instruction manual can yield hundreds of evaluation samples, enabling efficient benchmark expansion, while the availability of tens of thousands of open-source LEGO models across diverse categories ensures high scalability and diversity of the benchmark.
@@
Leveraging LEGO-Puzzles, we conduct comprehensive evaluations of 20 state-of-the-art MLLMs, including proprietary models such as GPT-4o and Gemini-2.0-Flash, as well as leading open-source alternatives [8, 49, 55, 42]. Our experimental results reveal a substantial gap between current MLLMs and human-level proficiency. Even the strongest models struggle with basic spatial understanding tasks, such as accurately identifying the height relationships in 3D space. Among open-source models, only a few achieve performance notably above random guessing across different tasks.
@@
To further investigate whether MLLMs can transfer there spatial understanding and reasoning abilities to image generation, we design generation tasks based on LEGO-Puzzles. For instance, given an assembly illustration, an MLLM is tasked with generating an image of the intermediate state following the specified assembly operation. In these generation tests, most of the evaluated models fail completely, either disregarding the provided instructions or generating images that are entirely irrelevant to the intended LEGO configuration.
@@
In summary, LEGO-Puzzles provides a comprehensive evaluation of the spatial understanding and sequential reasoning capabilities of MLLMs. Our main contributions are as follows:
@@
  • Evaluation for multi-step spatial reasoning. Built upon LEGO’s step-by-step building process, LEGO-Puzzles is the first benchmark explicitly designed to assess multi-step spatial reasoning, where each task requires reasoning over up to 7 LEGO construction steps.
  • Progressive and comprehensive task coverage. Our benchmark includes a diverse set of tasks spanning basic spatial understanding, single-step reasoning, and multi-step reasoning. This enables systematic evaluation of MLLMs’ reasoning capabilities across increasing levels of spatial and sequential complexity.
  • Exploratory evaluation of reasoning transfer to image generation. Building upon LEGO-Puzzles, we further conduct exploratory experiments to examine whether MLLMs can transfer their spatial understanding and reasoning abilities to image generation tasks. This sheds light on the models’ capacity for multimodal generalization beyond recognition and question answering.
@@ 3 LEGO-Puzzles
In this section, we introduce LEGO-Puzzles, a diverse and comprehensive benchmark designed to evaluate the multi-step spatial reasoning capability of MLLMs in detail. Specifically, we first introduce the motivation and definition of each task in Sec. 3.1. Then, we introduce our dataset curation process, including data collection, question-answer generation, and quality control, in Sec. 3.2.
@@ 3.1 Task Definition
To enable a more comprehensive and progressively structured evaluation of multi-step spatial reasoning in MLLMs, we define three categories of tasks. This framework is grounded in insights from cognitive psychology and human developmental stages in acquiring spatial intelligence [39, 5, 54]. Using LEGO building as a concrete and intuitive example, we observe that humans typically develop spatial reasoning abilities in stages—from basic spatial understanding, to reasoning through individual assembly steps, and ultimately to reasoning across multiple sequential steps. Based on this developmental trajectory, our benchmark is divided into three categories, as illustrated in Fig. 2 (a).
Type 1: Spatial Understanding. This category focuses on the ability to understand the spatial relationships between each LEGO piece and how these pieces relate from different perspectives in 3D space: (1) Height: Distinguish the relative heights of LEGO objects. (2) Adjacency: Determine whether LEGO objects are adjacent or separated. (3) Rotation: Calculate the angle of rotation between a LEGO object and its corresponding rotated version. (4) Multiview: Predict the current LEGO status from different viewpoints.
Type 2: Single-Step Sequential Reasoning. Building upon spatial understanding, this category evaluates the model’s ability to reason through the dependencies and assembly logic. Tasks are designed based on single-step actions, mirroring how humans typically progress from spatial perception to carrying out one assembly step at a time during the building process: (5) Rotation Status: Determine whether a LEGO piece requires rotation before installation. In contrast to Task 3, this task focuses on reasoning from an assembly perspective, with finer granularity. (6) Position: Identify the correct assembly position to place the next LEGO piece. (7) Next-Step: Predict the next LEGO status based on the current status and the new pieces. (8) Dependency: Identify which pieces are necessary to transition from the current to the next assembly stage.
Type 3: Multi-Step Sequential Reasoning. The final category assesses the ability to reason across multiple steps in the assembly process. These tasks build upon the single-step reasoning skills from Type 2 and require planning over extended sequences (involving up to 7 intermediate stages): (9) Backwards: Identify the correct intermediate stage of a LEGO build from the full assembly state. (10) Ordering: Determine the correct assembly order of the provided final LEGO images. (11) Outlier: Detect the LEGO status that does not belong to the provided assembly sequence.
Figure 2: Problem statistics and data curation in LEGO-Puzzles.
In conclusion, LEGO-Puzzles consists of over 1,100 visual question-answering (VQA) pairs derived from 407 LEGO building instructions, encompassing 11 tasks across spatial understanding, single-step and multi-step sequential reasoning. Each task contains 100 samples to ensure balanced evaluation across categories.
@@ 3.2 Dataset Curation
As illustrated in Fig. 2, our pipeline consists of three key steps: data collection, question-answer generation, and quality control. This design ensures the scalability, accuracy, and reliability of our data.
Data Collection. Data collection consists of three stages. (1). LEGO Project Collection. We collect a diverse set of open-source LEGO source files from the Internet, each containing detailed step-by-step building instructions and part lists. To ensure suitable task complexity, we filter for projects with moderate final size. Extremely large builds exhibit high structural complexity, and the small visual changes introduced by added pieces make it difficult for models to detect step-wise differences. Conversely, overly small builds are filtered out for low spatial complexity and insufficient steps for multi-step reasoning. We also ensure category diversity, covering animals, furniture, vehicles, and more, to increase task variety. (2). Rendering and Transformation. We render each project into PDF format using the publicly available software Studio [^1] https://www.bricklink.com/v3/studio/download.page, keeping the camera viewpoint fixed across steps to maintain spatial and temporal consistency. The tool allows flexible editing of the source files, enabling us to modify part attributes such as type, quantity, color, and position as needed for task construction. For example, in the Rotation and Multiview task, we apply POV-Ray style rendering and adjust lighting to simulate different viewing angles. In the Backward task, we introduce deliberate errors in part attributes to generate incorrect assembly states. (3).Metadata Extraction and Unification. We employ PDF-Extract [^2] https://github.com/opendatalab/PDF-Extract-Kit to extract structured information from the rendered PDFs, including individual LEGO pieces and assembled objects. All visual assets are processed under a unified naming convention and organized for downstream question-answer generation across all defined task types.
Question-Answer Generation. To support scalable and structured data construction, we manually design task-specific templates tailored to each task. Each example includes an instruction defining the model’s role, a question referencing input images using tokens like <image x>, and a ground-truth answer. For example, in the Position task, the model is provided with the current assembly state, the part to install, the resulting state, and several candidate placements. This standardized template design allows flexibility in the number of input images required by each task, making it suitable for both single-step and multi-step reasoning scenarios.
Quality Control. To ensure data quality and reliability, we implement a multi-stage human-in-the-loop review process. 1)Duplicate filtering. We perform duplication checks by computing similarity scores across rendered images, identifying visually redundant or recolored samples. These are further reviewed manually, and duplicates are removed to reduce redundancy and improve evaluation quality. 2)Image Quality and Format Check. We manually verify all rendered images to ensure consistency in camera perspective, correctness of part attributes (e.g., color, shape, quantity), and adherence to naming conventions. Any files with errors are either corrected or discarded. 3)Template and annotation validation. Each question-answer pair is verified by three trained annotators. Reviewers confirm that the images referenced by tokens (<image 1>, <image 2>, …) appear in the correct order. Samples with unresolved disagreements are either revised or removed.
Given the vast number of high-quality open-source LEGO projects available, our pipeline is inherently scalable and can be extended in an semi-automated manner to support larger and more diverse benchmarks in the future.
4 Experiment
@@ 4.1 Experimental Setting
Benchmark Models. We extensively evaluate 20 models, covering a diverse range of architectures, sizes, and training processes for Spatial Understanding and Sequential Reasoning tasks. For open-source models, we evaluate MiniCPM-V2.6 [57], Qwen2-VL-[7B/72B] [49], Qwen2.5-VL-[7B/72B] [4], InternVL2.5-[8B/78B] [7], VILA1.5-13B [31], Idefics3-8B [23], DeepSeek-VL2-[Tiny/Small] [55], Pixtral-12B [1], LLaVA-OneVision-7B [25], and EMU3 [53]. For proprietary models, we evaluate Claude-3.5-Sonnet [2], Gemini-1.5-Flash, Gemini-1.5-Pro, Gemini-2.0-Flash [47], GPT-4o (20241120), and GPT-4o-mini [41]. Moreover, all evaluations are conducted in a zero-shot setting for a fair comparison.
@@
Evaluation Metrics. For all questions in LEGO-Puzzles, we adopt accuracy (%) as the evaluation metric. For multiple-choice questions, we follow VLMEvalKit [11], applying rule-based matching as a first step, and resort to LLM-based choice matching using ChatGPT [41] when heuristic matching fails. For the Ordering task, we adopt the same strategy: rule-based extraction of the predicted step sequence from the model’s response, followed by LLM-based matching when necessary.
Table 1: Full evaluation results of 20 MLLMs on LEGO-Puzzles. Dark Gray and Light Gray indicates the best performance for each task among all models and open-source models respectively. We highlight the top 3 models by overall performace using Dark Green, Medium Green, and Light Green, respectively.
@@
Table 2: Comparing top-performing MLLMs with human proficiency on LEGO-Puzzles-Lite. The best results are marked in bold. The top 3 overall performances are highlighted in Dark Green, Medium Green, and Light Green, respectively.